Circular No. 6398

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 J1 (EVANS-DRINKWATER) Further precise positions have been reported as follows:

1996 UT α_{2000} δ_{2000} m_{1} Observer
May 10.44268 10^{h}52^{m}55.^{s}94 −23^{o}22′52′′.0 16  Drinkwater
12.60872 10 52 35.92 −23 06 05.6 16.5 Garradd
12.61565 10 52 35.95 −23 06 02.5  _{”}
13.11309 10 52 32.44 −23 02 17.2 15.8 Offutt
13.21946 10 52 31.59 −23 01 27.5 15.8  _{”}
13.44792 10 52 30.21 −22 59 43.4 15.6 Kushida
13.45903 10 52 30.11 −22 59 37.0  _{”}
13.47424 10 52 29.95 −22 59 27.5  Garradd
13.47698 10 52 29.93 −22 59 26.6  _{”}

M. J. Drinkwater (Siding Spring). Measurer R. H. McNaught. G. J. Garradd (Loomberah). 0.25-m f/4.1 reflector + CCD. W. Offutt (Cloudcroft). 0.60-m f/7 Ritchey-Chrétien + CCD. Y. Kushida (Yatsugatake). 0.40-m f/5 reflector + CCD. Communicated

by S. Nakano. Parabolic orbital elements from 11 observations May 10–13:

T = 1996 Dec. 27.857 TT  ω = 13.^{o}763
Ω = 278.908 }2000.0
q = 1.30192 AU  i = 22.457
1996TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
May 12 10^{h}52.^{m}68 −23^{o}10.^{′}8 2.708 3.312 118.^{o}5 15.^{o}5 15.9
17 10 52.26 −22 32.9 2.712 3.260 114.2 16.4 15.8
22 10 52.42 −21 56.8 2.719 3.208 109.9 17.3 15.7
27 10 53.15 −21 23.1 2.727 3.155 105.7 18.0 15.7
June 1 10 54.44 −20 52.0 2.737 3.103 101.5 18.7 15.6
6 10 56.28 −20 24.0 2.748 3.050  97.4 19.3 15.5

R. M. West and L. Ramirez, European Southern Observatory, have obtained four 5-min R frames of comet C/1996 J1 with the MPI/ESO 2.2m telescope + EFOSC II at La Silla at May 13.15 UT. The outer contours of the coma are elliptical with largest diameters 90′′ and 120′′ in p.a. 45^{o} and 315^{o}, respectively, and integrated R = 13. There is a broad tail in p.a. 315^{o}. The central light peak is diffuse and is surrounded by a fanlike structure opening over 90^{o} symmetrically around p.a. 0^{o}.

1996 May 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Brian G. Marsden